The distribution of iodine between carbon tetrachloride and water is constant. The value of K is cccia / cH2o = 85. How many ml of carbon tetrachloride would be required to extract 959% iodine from 100 ml of aqueous solution in a single phase? 


1
Expert's answer
2021-04-02T05:12:26-0400

"K = \dfrac{C_{CCl_4}}{C_{H_2O}} = 85"

Suppose we have 1 mmol / L solution of I2 in water. Then 0.959 mmol must be extracted, and 0.041 mmol left in water.

"C = \dfrac{n}{V}"

C(iodine in water) = 0.041 / 0.1 = 0.41 mmol / L

85 = "\frac{X}{0.41}", x = "0.41 \times 85" = 34.85 (mmol / L)

VCCl4 = 0.959 / 34.85 = 0.0275 L = 27.5 ml


Answer: 27.5 ml.